ganeshie8 (ganeshie8):

What is the best way to simplify 2103/2127

ganeshie8 (ganeshie8):

\[\dfrac{2103}{2127}\]

OpenStudy (kmeezy):

simplified is 701/709 by finding out the common factor which is 3

ganeshie8 (ganeshie8):

Nice! how do we know 701 and 709 have no common factors ?

OpenStudy (astrophysics):

Prime numbers

ganeshie8 (ganeshie8):

Is it easy to figure that out

OpenStudy (samigupta8):

We can write it as 1-8/709

OpenStudy (samigupta8):

So no common factors !

ganeshie8 (ganeshie8):

Interesting... could you explain how you conclude no common factors from that ?

OpenStudy (samigupta8):

Since 8 and 709 don't have any factors in common, i concluded that.

OpenStudy (samigupta8):

Like say we have this 16/4 as the number . If i write it to be 12+4 /4 then i get 1+12/4 Here 4 and 12 have common thing in them .So we can accordingly get the factors.

OpenStudy (excalibur0126):

Here's a good website describing the rules of divisibility. http://usablealgebra.landmark.edu/algebra/factoring/divisibility-rules.php

OpenStudy (baru):

i agree with @samigupta8 consider the mixed fraction \[1-\frac{a}{b}\] such that a and b have no common factors. then \[1-\frac{a}{b}= \frac{b-a}{b}\] let if (b-a) and (a) have common factor (k), we can rewrite as \[\frac{b-a}{b}=\frac{nk}{mk} \\or\\b-a=nk\\b=mk\] subtract the above two equaitons we get a=(m-n)k but we also have b=mk so a and b have common factor 'k' which is contradiction with our initial suppositon that a and b have no common factors

ganeshie8 (ganeshie8):

Awesome! really a great observation and a cute proof !

OpenStudy (baru):

yay :)

ganeshie8 (ganeshie8):

Loks you have derived the euclidean algorithm \[\gcd(b,~a) = \gcd(b, ~b-a)\]

ganeshie8 (ganeshie8):

@baru @samigupta8